f(x)=2x + 1 and g(x)=2x(squared) + x
Determine the following:
#13 (f of g)(2)

Answer by Nate(3500)

f(x)=2x + 1 and g(x)=2x(squared) + x
[f o g](2) plug g(x) into f(x)
f(x)=2x + 1
f(x)=2(2x^2 + x) + 1
f(x)=4x^2 + 2x + 1
f(2)=4(2^2) + 2(2) + 1
f(2)=16 + 4 + 1
f(2)=21